# FedSA-GCL: A Semi-Asynchronous Federated Graph Learning Framework with Personalized Aggregation and Cluster-Aware Broadcasting

### v1

@@ Abstract
Federated Graph Learning (FGL) is a distributed learning paradigm that enables collaborative training over large-scale subgraphs located on multiple local systems. However, most existing FGL approaches rely on synchronous communication, which leads to inefficiencies and is often impractical in real-world deployments. Meanwhile, current asynchronous federated learning (AFL) methods are primarily designed for conventional tasks such as image classification and natural language processing, without accounting for the unique topological properties of graph data. Directly applying these methods to graph learning can possibly result in semantic drift and representational inconsistency in the global model. To address these challenges, we propose FedSA-GCL, a semi-asynchronous federated framework that leverages both inter-client label distribution divergence and graph topological characteristics through a novel ClusterCast mechanism for efficient training. We evaluate FedSA-GCL on multiple real-world graph datasets using the Louvain and Metis split algorithm, and compare it against 9 baselines. Extensive experiments demonstrate that our method achieves strong robustness and outstanding efficiency, outperforming the baselines by an average of 2.92% with the Louvain and by 3.4% with the Metis.
@@
journal: Expert Systems with Applications
@@
Achieves 484.5 fewer trips over baseline, improving communication efficiency.
@@ 1 Introduction
Graphs are data structure for modeling complex relational systems across a wide range of domains, including social networks(Guo & Wang, 2020), drug discovery(Liu et al., 2022)(Wang et al., 2023), and recommendation platforms(He et al., 2023)(Yu et al., 2022). Although graph machine learning has made significant progress in recent years, most existing methods rely on centralized storage of large-scale graph data on a single server. However, with growing concerns about data security and user privacy, this centralized model is often impractical in real-world scenarios. In practice, graph data is typically distributed across multiple data owners, and cross-site data collection becomes impractical due to privacy restrictions(Fu et al., 2022).
@@
To address this challenge, the concept of Federated Graph Learning (FGL) has emerged, which integrates federated learning (FL) with graph learning. FGL enables collaborative model training among distributed graph data holders without requiring raw data exchange, thereby preserving privacy. In FGL, graph data is typically partitioned into local subgraphs held by different clients, such as transportation subnet works (Zhang et al., 2021b; Li & Zhu, 2021), social communities (Guo & Wang, 2020), or financial chains (Yang et al., 2021). A key challenge for FGL is subgraph heterogeneity, which refers to the significant divergence in label distributions and graph structures—driven by strong node dependencies—across clients(Baek et al., 2023; Xie et al., 2021), thereby undermining cross-client model generalization. To mitigate the impact of subgraph heterogeneity, recent studies in FGL have begun to explicitly account for the unique structural and distributional characteristics of graph data. For example, several works incorporate topology-aware designs to capture local graph semantics (Zhang et al., 2021a; Li et al., 2024b), while others explore personalized modeling or representation-level adaptation to better handle client-specific variations (Tan et al., 2022; Baek et al., 2023; Li et al., 2024a).
@@
Despite their effectiveness, these methods are still primarily designed under synchronous training protocols, which pose significant bottlenecks to efficiency in practical settings (Jiang et al., 2022). Consequently, current FGL research continues to face two major limitations.
@@
L1: Synchronous FGL Suffers from Inefficiency. Although existing FGL methods have demonstrated promising results (Baek et al., 2023; Li et al., 2024b, a; Zhu et al., 2024), they are predominantly designed for idealized synchronous settings. However, their reliance on synchronization raises serious concerns for practical deployment. In heterogeneous environments, such as those involving edge devices or clients with limited computing capacity, stragglers can delay the entire training process, leading to prolonged communication rounds and underutilized system resources (Chen et al., 2019). Additionally, in real-world scenarios, client participation is often sporadic and inconsistent. Strictly requiring all selected clients to complete training before aggregation is incompatible with such dynamic participation patterns, rendering synchronous protocols not only inefficient but often infeasible.
@@
L2: Asynchronous Settings Amplify Graph Non-IID. Existing asynchronous federated learning (AFL) methods (Xie et al., 2019; Xu et al., 2023; Ma et al., 2021; Nguyen et al., 2022) are primarily developed for domains such as computer vision and natural language processing. These approaches typically assume independent and identically distributed (IID) data, and lack dedicated mechanisms to address the cross-node dependencies and topological heterogeneity intrinsic to graph data. Although SWESALT (Liao et al., 2023) introduces asynchronous communication into FGL, it does not explicitly address the Non-IID nature of graph-structured data. The inherent non-IID nature of graph data arises from topological heterogeneity and cross-node dependencies. While synchronous FL can mitigate some of this heterogeneity by aggregating updates from all clients simultaneously, asynchronous participation disrupts this balance. The server often aggregates updates from a temporally sparse and biased subset of clients, causing certain structural patterns to be underrepresented. This amplifies the existing non-IID challenge, further degrading the consistency and generalizability of the global model.
@@ 2.1 Problem Formulation
Due to the lack of publicly available benchmark datasets for federated graph learning (FGL), existing works (Zhang et al., 2021a; Chen et al., 2024; Baek et al., 2023) typically simulate FGL scenarios by splitting a single global graph into multiple subgraphs. Given a global graph $G=(\mathcal{V},\mathcal{E})$ with $|\mathcal{V}|=n$ nodes and $|\mathcal{E}|=m$ edges, its structure can be represented by an adjacency matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$. Each node $v\in\mathcal{V}$ is associated with a feature vector $\mathbf{x}_{v}\in\mathbb{R}^{f}$, forming the feature matrix $\mathbf{X}\in\mathbb{R}^{n\times f}$, and a one-hot label vector $\mathbf{y}_{v}\in\mathbb{R}^{|\mathcal{Y}|}$, resulting in the label matrix $\mathbf{Y}\in\mathbb{R}^{n\times|\mathcal{Y}|}$, where $|\mathcal{Y}|$ is the number of classes. To simulate a federated environment, community detection algorithms such as Metis(Karypis & Kumar, 1998) or Louvain(Blondel et al., 2008) are applied to split the global graph $G$ into $N$ subgraphs. Each client $i\in\left\{1,\dots,N\right\}$ is assigned a local subgraph $G_{i}=(\mathcal{V}_{i},\mathcal{E}_{i})$, along with the corresponding local adjacency matrix $\mathbf{A}_{i}\in\mathbb{R}^{|\mathcal{V}_{i}|\times|\mathcal{V}_{i}|}$, feature matrix $\mathbf{X}_{i}\in\mathbb{R}^{|\mathcal{V}_{i}|\times f}$, and label matrix $\mathbf{Y}_{i}\in\mathbb{R}^{|\mathcal{V}_{i}|\times|\mathcal{Y}|}$.
@@ 2.2 Federated Graph Learning
FedAvg (McMahan et al., 2017) is the most widely used optimization strategy in federated learning. It performs global model aggregation by computing a simple weighted average of the model parameters received from participating clients, where the weights are typically proportional to the size of local training data. In this setting, we integrate graph neural networks (GNNs)(Kipf & Welling, 2016) with FedAvg. The aggregation process at round $t$ is given by Eq. 1.
@@
Although FedAvg is widely adopted, it performs poorly under non-IID client data and fails to capture the unique properties of graph-structured data. To address this, FedSage+(Zhang et al., 2021a) integrates GraphSAGE into the federated setting and reconstructs missing inter-subgraph edges, but mainly focuses on structural completion. FedProto(Tan et al., 2022) tackles label distribution heterogeneity via prototype sharing, yet ignores graph topology. FedPub (Baek et al., 2023) introduces functional embeddings and sparse aggregation to enable personalized training on heterogeneous subgraphs. FedGTA (Li et al., 2024b) and AdaFGL (Li et al., 2024a) incorporate topological confidence and structure-aware personalization strategies, improving both convergence and adaptability. More recently, FedTAD (Zhu et al., 2024) adopts topology-aware, data-free distillation to enable structure-sensitive aggregation without data sharing.
@@ 2.3 Asynchronous Federated Learning
To overcome the inefficiencies of synchronous federated learning under system heterogeneity, Asynchronous Federated Learning (AFL) has been proposed to enable non-blocking client-server communication (Xu et al., 2023). AFL can be broadly categorized based on its aggregation strategy into fully asynchronous and semi-asynchronous approaches. Fully asynchronous methods such as FedAsync (Xie et al., 2019) and TWAFL (Chen et al., 2019) allow the server to update the global model immediately upon receiving a single client update, improving reactivity but risking instability under non-IID conditions. In contrast, semi-asynchronous methods like FedSA (Ma et al., 2021) and FedBuff (Nguyen et al., 2022) trigger aggregation only after collecting a certain number of updates, striking a balance between convergence stability and communication efficiency. TimelyFL (Zhang et al., 2023) introduces a tier-based scheduling mechanism that adapts local training based on device availability and latency, further enhancing responsiveness in heterogeneous environments.
Although some preliminary efforts, such as SWESALT (Liao et al., 2023), have extended semi-asynchronous strategies to graph data, they do not incorporate graph-specific optimizations and overlook the non-IID nature of graph data. Current AFL methods typically assume IID input and lack mechanisms to address graph-specific challenges such as topological heterogeneity and cross-node dependencies. As a result, their direct application to graph learning often leads to unstable convergence and limited generalization, highlighting the need for graph-aware asynchronous federated algorithms.
@@ 3.1 Client Clustering Based on Soft Label Feature Matrix
where $G_{i}(\mathcal{V}_{i},\mathcal{E}_{i})$ denotes the local subgraph of client i, and $f(\tilde{\omega_{i}};\cdot)$is the predictive model parameterized by the aggregated weights $\tilde{\omega}_{i}$.
@@ 3.2 Model Aggregation Based on Local Smoothness Confidence and Staleness
To quantify the topological-attribute correlation between nodes, we introduce k-step Non-parametric Label Propagation (Non-param LP), which models the relationship between target nodes and their k-hop neighbors through iterative updates(Gasteiger et al., 2018).To capture higher-order structural information, the $k$-step label propagation is given in Eq. (6).
@@
The entropy of node soft labels quantifies the prediction consistency of local models, while the local smoothing confidence characterizes their consistency with neighborhood predictions(Li et al., 2024b). Specifically: Lower entropy indicates stronger prediction consistency within local graph structures, yielding more reliable learned topological features. Furthermore, considering the significant influence of high-degree nodes on local smoothing confidence, we assign them higher weights. The local smoothness confidence (LSC), which reflects the entropy-based consistency between nodes and their neighbors, as shown in Eq. (7).
@@
In asynchronous federated learning frameworks, clients conduct local training based on different versions of the global model, resulting in model version staleness(Xie et al., 2019). To address this issue, we employ a staleness metric to quantify version discrepancy and accordingly adjust client model aggregation weights. The staleness-aware aggregation weight is given in Eq. (8):
@@ 3.3 Semi-Asynchronous ClusterCast Broadcasting
A key challenge in semi-asynchronous federated learning is maintaining a timely and coherent information flow under partial client participation. Existing methods like FedBuff(Nguyen et al., 2022) trigger aggregation upon receiving a fixed number of client updates, but only send the aggregated model to those clients who participated in that current round. As illustrated in Fig. LABEL:fig:comm_diagram_sub1, this limitation can cause structurally similar but inactive clients—such as client 3 in round 2—to miss out on valuable updates, ultimately leading to slower convergence and increased inter-client drift. To address this issue, we propose ClusterCast, a novel broadcasting mechanism designed to enhance cluster-level cross-client consistency and training efficiency.
@@
1:Initialize: $t\leftarrow 0$, $\mathcal{U}\leftarrow\emptyset$, $\mathcal{B}\leftarrow\emptyset$triangleright Global round, uploaded client set, broadcast set
@@
[…]
4:     $t\leftarrow t+1$,$\mathcal{U}\leftarrow\emptyset$$\mathcal{B}\leftarrow\emptyset$
[…]
@@
3:     Receive $(\tilde{\omega},\tau,\tilde{LSC})$ from DownloadQueue.get()
4:     Replace stored update with $(\tilde{\omega},\tau,\tilde{LSC})$triangleright Keep only the latest update
@@ 4.1 Experiments Setup
Dataset. To comprehensively evaluate the performance of our method against other baselines, we select seven graph datasets. Our experiments are conducted on three small-scale citation networks (Cora, CiteSeer, PubMed) (Yang et al., 2016), two medium-scale user-item datasets (Amazon Computers, Amazon Photo), and two medium-scale co-authorship networks (Coauthor CS, Coauthor Physics) (Shchur et al., 2018). Detailed descriptions of these datasets and the proportions of the train/validation/test splits are presented in Table 1.
To simulate realistic federated graph learning scenarios, we apply two community detection algorithms, Metis(Karypis & Kumar, 1998) and Louvain(Blondel et al., 2008), to split each dataset into 20 clients. These splits are used for federated graph training. For several representative datasets and both community detection algorithms, the label distributions across clients are illustrated in Fig. 3.
@@
Baselines. We compare our method against nine baselines. For traditional federated learning (FL), we select FedAvg(McMahan et al., 2017), FedProx(Li et al., 2020), and MOON(Li et al., 2021). For federated graph learning (FGL), we include FedSage+(Zhang et al., 2021a), FedGTA(Li et al., 2024b), and FedTAD(Zhu et al., 2024). For asynchronous federated learning (AFL), we compare with FedAnsy(Xie et al., 2019), TWAFL(Chen et al., 2019), and FedBuff(Nguyen et al., 2022). All experiments are conducted using the GCN model as the backbone architecture.
Hyperparameters. Our asynchronous training scripts are developed based on the open-source federated graph learning framework OpenFGL(Li et al., 2024c). For baseline methods, we adopt the hyperparameter settings reported in their original papers. In federated learning, the number of communication rounds to reach a target accuracy is a common evaluation metric(Nguyen et al., 2022). However, since asynchronous and synchronous approaches differ in their notion of a round, FedBuff proposes the client trip — a cycle of model download, local training for one epoch, and model upload—as a fair basis for comparison. Following this practice, we also adopt the client trip as the evaluation unit to ensure fair comparison across methods. We set the number of local epochs to 1 and train for 2000 client trip. To simulate an asynchronous environment, 30% of clients are randomly designated as edge devices, which fall behind the global server by 2 to 5 complete training cycles (e.g., with 20 clients, one global training cycle corresponds to 20 client trip). For cluster-based federated learning, the clustering threshold is set to 0.5, as suggested in FedGTA. Model validation is performed locally on each client’s local dataset.

### v2

@@ Abstract
Federated Graph Learning (FGL) is a distributed learning paradigm that enables collaborative training over large-scale subgraphs located on multiple local systems. However, most existing FGL approaches rely on synchronous communication, which leads to inefficiencies and is often impractical in real-world deployments. Meanwhile, current asynchronous federated learning (AFL) methods are primarily designed for conventional tasks such as image classification and natural language processing, without accounting for the unique topological properties of graph data. Directly applying these methods to graph learning can possibly result in semantic drift and representational inconsistency in the global model. To address these challenges, we propose FedSA-GCL, a semi-asynchronous federated framework that leverages both inter-client label distribution divergence and graph topological characteristics through a novel ClusterCast mechanism for efficient training. We evaluate FedSA-GCL on multiple real-world graph datasets using the Louvain and Metis algorithms, and compare it against 9 baselines. Extensive experiments demonstrate that our method achieves strong robustness and outstanding efficiency, outperforming the baselines by an average of 2.92% with the Louvain and by 3.4% with the Metis.
@@
journal:
@@
Reduces client trips by 484.5 over baseline, improving communication efficiency.
@@ 1 Introduction
Graphs are data structure for modeling complex relational systems across a wide range of domains, including social networks[1], drug discovery[2, 3], and recommendation platforms[4, 5]. Although graph machine learning has made significant progress in recent years, most existing methods rely on centralized storage of large-scale graph data on a single server. However, with growing concerns about data security and user privacy, this centralized model is often impractical in real-world scenarios. In practice, graph data is typically distributed across multiple data owners, and cross-site data collection becomes impractical due to privacy restrictions[6].
@@
To address this challenge, the concept of Federated Graph Learning (FGL) has emerged, which integrates federated learning (FL) with graph learning. FGL enables collaborative model training among distributed graph data holders without requiring raw data exchange, thereby preserving privacy. In FGL, graph data is typically partitioned into local subgraphs held by different clients, such as transportation subnet works [7, 8], social communities [1], or financial chains [9]. A key challenge for FGL is subgraph heterogeneity, which refers to the significant divergence in label distributions and graph structures—driven by strong node dependencies—across clients[10, 11], thereby undermining cross-client model generalization. To mitigate the impact of subgraph heterogeneity, recent studies in FGL have begun to explicitly account for the unique structural and distributional characteristics of graph data. For example, several works incorporate topology-aware designs to capture local graph semantics [12, 13], while others explore personalized modeling or representation-level adaptation to better handle client-specific variations [14, 10, 15].
@@
Despite their effectiveness, these methods are still primarily designed under synchronous training protocols, which pose significant bottlenecks to efficiency in practical settings [16]. Consequently, current FGL research continues to face two major limitations.
@@
L1: Synchronous FGL Suffers from Inefficiency. Although existing FGL methods have demonstrated promising results [10, 13, 15, 17], they are predominantly designed for idealized synchronous settings. However, their reliance on synchronization raises serious concerns for practical deployment. In heterogeneous environments, such as those involving edge devices or clients with limited computing capacity, stragglers can delay the entire training process, leading to prolonged communication rounds and underutilized system resources [18]. Additionally, in real-world scenarios, client participation is often sporadic and inconsistent. Strictly requiring all selected clients to complete training before aggregation is incompatible with such dynamic participation patterns, rendering synchronous protocols not only inefficient but often infeasible.
@@
L2: Asynchronous Settings Amplify Graph Non-IID. Existing asynchronous federated learning (AFL) methods [19, 20, 21, 22] are primarily developed for domains such as computer vision and natural language processing. These approaches typically assume independent and identically distributed (IID) data, and lack dedicated mechanisms to address the cross-node dependencies and topological heterogeneity intrinsic to graph data. Although SWESALT [23] introduces asynchronous communication into FGL, it does not explicitly address the Non-IID nature of graph-structured data. The inherent non-IID nature of graph data arises from topological heterogeneity and cross-node dependencies. While synchronous FL can mitigate some of this heterogeneity by aggregating updates from all clients simultaneously, asynchronous participation disrupts this balance. The server often aggregates updates from a temporally sparse and biased subset of clients, causing certain structural patterns to be underrepresented. This amplifies the existing non-IID challenge, further degrading the consistency and generalizability of the global model.
@@ 2.1 Problem Formulation
Due to the lack of publicly available benchmark datasets for federated graph learning (FGL), existing works [12, 24, 10] typically simulate FGL scenarios by splitting a single global graph into multiple subgraphs. Given a global graph $G=(\mathcal{V},\mathcal{E})$ with $|\mathcal{V}|=n$ nodes and $|\mathcal{E}|=m$ edges, its structure can be represented by an adjacency matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$. Each node $v\in\mathcal{V}$ is associated with a feature vector $\mathbf{x}_{v}\in\mathbb{R}^{f}$, forming the feature matrix $\mathbf{X}\in\mathbb{R}^{n\times f}$, and a one-hot label vector $\mathbf{y}_{v}\in\mathbb{R}^{|\mathcal{Y}|}$, resulting in the label matrix $\mathbf{Y}\in\mathbb{R}^{n\times|\mathcal{Y}|}$, where $|\mathcal{Y}|$ is the number of classes. To simulate a federated environment, community detection algorithms such as Metis[25] or Louvain[26] are applied to split the global graph $G$ into $N$ subgraphs. Each client $i\in\left\{1,\dots,N\right\}$ is assigned a local subgraph $G_{i}=(\mathcal{V}_{i},\mathcal{E}_{i})$, along with the corresponding local adjacency matrix $\mathbf{A}_{i}\in\mathbb{R}^{|\mathcal{V}_{i}|\times|\mathcal{V}_{i}|}$, feature matrix $\mathbf{X}_{i}\in\mathbb{R}^{|\mathcal{V}_{i}|\times f}$, and label matrix $\mathbf{Y}_{i}\in\mathbb{R}^{|\mathcal{V}_{i}|\times|\mathcal{Y}|}$.
@@ 2.2 Federated Graph Learning
FedAvg [27] is the most widely used optimization strategy in federated learning. It performs global model aggregation by computing a simple weighted average of the model parameters received from participating clients, where the weights are typically proportional to the size of local training data. In this setting, we integrate graph neural networks (GNNs)[28] with FedAvg. The aggregation process at round $t$ is given by Eq. 1.
@@
Although FedAvg is widely adopted, it performs poorly under non-IID client data and fails to capture the unique properties of graph-structured data. To address this, FedSage+[12] integrates GraphSAGE into the federated setting and reconstructs missing inter-subgraph edges, but mainly focuses on structural completion. FedProto[14] tackles label distribution heterogeneity via prototype sharing, yet ignores graph topology. FedPub [10] introduces functional embeddings and sparse aggregation to enable personalized training on heterogeneous subgraphs. FedGTA [13] and AdaFGL [15] incorporate topological confidence and structure-aware personalization strategies, improving both convergence and adaptability. More recently, FedTAD [17] adopts topology-aware, data-free distillation to enable structure-sensitive aggregation without data sharing.
@@ 2.3 Asynchronous Federated Learning
To overcome the inefficiencies of synchronous federated learning under system heterogeneity, Asynchronous Federated Learning (AFL) has been proposed to enable non-blocking client-server communication [20]. AFL can be broadly categorized based on its aggregation strategy into fully asynchronous and semi-asynchronous approaches. Fully asynchronous methods such as FedAsync [19] and TWAFL [18] allow the server to update the global model immediately upon receiving a single client update, improving reactivity but risking instability under non-IID conditions. In contrast, semi-asynchronous methods like FedSA [21] and FedBuff [22] trigger aggregation only after collecting a certain number of updates, striking a balance between convergence stability and communication efficiency. TimelyFL [29] introduces a tier-based scheduling mechanism that adapts local training based on device availability and latency, further enhancing responsiveness in heterogeneous environments.
Although some preliminary efforts, such as SWESALT [23], have extended semi-asynchronous strategies to graph data, they do not incorporate graph-specific optimizations and overlook the non-IID nature of graph data. Current AFL methods typically assume IID input and lack mechanisms to address graph-specific challenges such as topological heterogeneity and cross-node dependencies. As a result, their direct application to graph learning often leads to unstable convergence and limited generalization, highlighting the need for graph-aware asynchronous federated algorithms.
@@ 3.1 Client Clustering Based on Soft Label Feature Matrix
[…]
@@ 3.2 Model Aggregation Based on Local Smoothness Confidence and Staleness
To quantify the topological-attribute correlation between nodes, we introduce k-step Non-parametric Label Propagation (Non-param LP), which models the relationship between target nodes and their k-hop neighbors through iterative updates[30].To capture higher-order structural information, the $k$-step label propagation is given in Eq. (6).
@@
The entropy of node soft labels quantifies the prediction consistency of local models, while the local smoothing confidence characterizes their consistency with neighborhood predictions[13]. Specifically: Lower entropy indicates stronger prediction consistency within local graph structures, yielding more reliable learned topological features. Furthermore, considering the significant influence of high-degree nodes on local smoothing confidence, we assign them higher weights. The local smoothness confidence (LSC), which reflects the entropy-based consistency between nodes and their neighbors, as shown in Eq. (7).
@@
In asynchronous federated learning frameworks, clients conduct local training based on different versions of the global model, resulting in model version staleness[19]. To address this issue, we employ a staleness metric to quantify version discrepancy and accordingly adjust client model aggregation weights. The staleness-aware aggregation weight is given in Eq. (8):
@@ 3.3 Semi-Asynchronous ClusterCast Broadcasting
A key challenge in semi-asynchronous federated learning is maintaining a timely and coherent information flow under partial client participation. Existing methods like FedBuff[22] trigger aggregation upon receiving a fixed number of client updates, but only send the aggregated model to those clients who participated in that current round. As illustrated in Fig. LABEL:fig:comm_diagram_sub1, this limitation can cause structurally similar but inactive clients—such as client 3 in round 2—to miss out on valuable updates, ultimately leading to slower convergence and increased inter-client drift. To address this issue, we propose ClusterCast, a novel broadcasting mechanism designed to enhance cluster-level cross-client consistency and training efficiency.
@@
1:Initialize: $t\leftarrow 0$, $\mathcal{U}\leftarrow\emptyset$, $\mathcal{B}\leftarrow\emptyset$ $\triangleright$ Global round, uploaded client set, broadcast set
@@
3:while $\text{UploadQueue.qsize()}\geq K$ do
4:  $t\leftarrow t+1$$\mathcal{U}\leftarrow\emptyset$$\mathcal{B}\leftarrow\emptyset$
5:  while $\text{UploadQueue.qsize()}\neq 0$ do
@@
1:Initialize: $\tau\leftarrow 0$, $\texttt{DownloadQueue}\leftarrow\texttt{Queue()}$,update $\leftarrow$ null $\triangleright$ Python thread-safe queue for receiving server updates
2:while $\texttt{DownloadQueue.qsize()}\neq 0$ do
@@ 4.1 Experiments Setup
Dataset. To comprehensively evaluate the performance of our method against other baselines, we select seven graph datasets. Our experiments are conducted on three small-scale citation networks (Cora, CiteSeer, PubMed) [31], two medium-scale user-item datasets (Amazon Computers, Amazon Photo), and two medium-scale co-authorship networks (Coauthor CS, Coauthor Physics) [32]. Detailed descriptions of these datasets and the proportions of the train/validation/test splits are presented in Table 1.
To simulate realistic federated graph learning scenarios, we apply two community detection algorithms, Metis[25] and Louvain[26], to split each dataset into 20 clients. These splits are used for federated graph training. For several representative datasets and both community detection algorithms, the label distributions across clients are illustrated in Fig. 3.
@@
Baselines. We compare our method against nine baselines. For traditional federated learning (FL), we select FedAvg[27], FedProx[33], and MOON[34]. For federated graph learning (FGL), we include FedSage+[12], FedGTA[13], and FedTAD[17]. For asynchronous federated learning (AFL), we compare with FedAnsy[19], TWAFL[18], and FedBuff[22]. All experiments are conducted using the GCN model as the backbone architecture.
Hyperparameters. Our asynchronous training scripts are developed based on the open-source federated graph learning framework OpenFGL[35]. For baseline methods, we adopt the hyperparameter settings reported in their original papers. In federated learning, the number of communication rounds to reach a target accuracy is a common evaluation metric[22]. However, since asynchronous and synchronous approaches differ in their notion of a round, FedBuff proposes the client trip — a cycle of model download, local training for one epoch, and model upload—as a fair basis for comparison. Following this practice, we also adopt the client trip as the evaluation unit to ensure fair comparison across methods. We set the number of local epochs to 1 and train for 2000 client trip. To simulate an asynchronous environment, 30% of clients are randomly designated as edge devices, which fall behind the global server by 2 to 5 complete training cycles (e.g., with 20 clients, one global training cycle corresponds to 20 client trip). For cluster-based federated learning, the clustering threshold is set to 0.5, as suggested in FedGTA. Model validation is performed locally on each client’s local dataset.